A line segment 42 inches long is divided into 3 portions in a ratio of 1:2:4. How many inches longer is the longest portion than the shortest portion?
A)16
B)18
C)20
D22

Answer by ikleyn(52793)

.
x + 2x + 4x = 42


7x = 42.


x = 42/7 = 6


ANSWER.  4x - x = 3x = 3*6 = 18 inches.